2y(t) + 3 y (t) + y(t)=x(t) +x(t) - x(t), y(0) = -2, y(0) = 0, u(t) is the step function. 1. Write an expression for Y(s
Answer #1

244) + 3 + 4 +44) = x^2+x4) X(4) 5- domain 2[5* Y29)- syro)?yko)] +3CsYess - yra] +Y44)= ($75 -1)x() (25 +3541) 968) + 4 = (